Parker Prints is in negotiation with two of its largest customers to increase the​ firm's sales dramatically. The increase will require that Parker expand its production facilities at a cost of $ 30 million. Parker expects to pay out $ 6 million in dividends to its shareholders next year. Parker maintains a 20 percent debt ratio in its capital structure.

a. If Parker earns $ 14 million next​ year, how much common stock will the firm need to sell in order to maintain its target capital​ structure?

b. If Parker wants to avoid selling any new​ stock, how much can the firm spend on new capital​ expenditures?
Answer #1

a) Retained Earnings for Parker after Dividend payout = 14 million - 6 million = 8 million
with 20% debt ratio amount of debt required = 20%*30 = 6 million
Extra equity required = Cost of Production facility - Retained Earnings - Debt required = 30 - 8 - 6 = 16 million

b) Retained earnings = 8 million
Debt ratio max = 20%
Debt Ratio = Debt/(Debt+8) = 20%
Debt = 2 million
Total Capital expenditure = Debt+ equity = 2+8 = 10 million

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well